Three point charges 10 C, -20 C, and -50 C are at the three consecutive corners of a square 10 cm on a side. What is the potential energy of the system?


1
Expert's answer
2022-03-14T13:06:56-0400

Potential energy

"U_1=\frac{kq_1q_2}{r}"

"U_1=\frac{k\times10\times(-20)}{0.1^2}"

"U_1=-{1.8\times10^{14}}J"

"U_2=\frac{kq_2q_3}{r^2}"

"U_2=\frac{k\times20\times50}{0.1^2}=9\times10^{14}J"

"U=\sqrt{U_1^2+U_2^2}"


"U=\sqrt{(-1.8\times10^{14})^2+(9\times10^{14})^2}=9.18\times10^{14}J"